OpenStudy (anonymous):

For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? -x + 2y = -6 3x + y = 8 -2y + 6 2y - 6 2y + 6 -2y - 6

OpenStudy (anonymous):

2y+6

OpenStudy (anonymous):

how'd you get that ?

OpenStudy (anonymous):

just -x+2y=-6 => 

OpenStudy (anonymous):

Oh okay, thank you!

OpenStudy (anonymous):

Can you help me with another one ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

pls hurry for that.

OpenStudy (anonymous):

A system of equations is shown below: 5x + 2y = 3 (equation 1) 2x – 3y = 1 (equation 2) A student wants to prove that if equation 2 is kept unchanged and equation 1 is replaced with the sum of equation 1 and a multiple of equation 2, the solution to the new system of equations is the same as the solution to the original system of equations. If equation 2 is multiplied by 1, which of the following steps should the student use for the proof? Show that the solution to the system of equations 7x – y = 4 and 2x – 3y = 1 is the same as the solution to the given system of equations Show that the solution to the system of equations 2x + 5y = 3 and 3x – 2y = 1 is the same as the solution to the given system of equations Show that the solution to the system of equations 9x + 4y = 5 and 7x – y = 4 is the same as the solution to the given system of equations Show that the solution to the system of equations –4x + 9y = 5 and 2x – 3y = 1 is the same as the solution to the given system of equations

OpenStudy (anonymous):

I'm not typing idk why it says I am.

OpenStudy (anonymous):

just wait a minute

OpenStudy (anonymous):

do u need any help

OpenStudy (anonymous):

i think it is 1st one since as per question on adding eq 1+eq 2 we get 7x-y=4

OpenStudy (anonymous):

and on multiplying eq. 2 by 1 we get the same 2x-3y=1 and on solving the above one and 7x-y=4 we get same solution as that from original equation give us solution

OpenStudy (anonymous):

hope u get that ? thanks for asking for help .

OpenStudy (anonymous):

hey r u there?